integrate the following integral
∫ (1/2-3x) + (1/sqrt(3x-2)) dx  

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int [\frac{1}{2-3x}+\frac{1}{\sqrt{3x-2}}]dx
I = \int \frac{1}{2-3x}dx+\int \frac{1}{\sqrt{3x-2}}dx
I = \frac{log(2-3x)}{-3}+\int \frac{1}{\sqrt{3x-2}}dx
I = -\frac{log(2-3x)}{3}+ \frac{2\sqrt{3x-2}}{3}+c